Emerald is 5 years older than Sapphire. Ruby is twice as old as Sapphire. The sum of the three ages is 65. How old is each?
Answer by mananth(16946)

Sapphire be x years
Emerald = x+5
Ruby = 2x
....
x+x+5+2x= 65
4x+5 = 65
4x = 60
x= 15 years sapphire's age
Emerald's age = x+5 = 15+5 = 20 years
Ruby's age = 2x = 2*15 = 30 years
...
m.ananth@hotmail.ca